### v1

# Integral representation of solutions to initial-boundary value problems in the framework of the hyperbolic heat equation

Abstract
We consider initial-boundary value problems on a finite interval for the system of the energy balance equation and modified Fourier’s law (constitutive equation, commonly called the Cattaneo equation) describing temperature (or internal energy) and heat flux. This hyperbolic system of first-order partial differential equations is the simplest model of non-Fourier heat conduction. Boundary conditions comprise various models of behavior of a physical system at the boundaries, including boundary conditions describing Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case the boundary conditions express the relationship of unknown functions (temperature or internal energy and heat flux) with each other. To solve the problems, we apply the Fokas unified transform method.
Keywords: non-Fourier heat conduction, Cattaneo equation, modified Fourier’s law, hyperbolic heat equation, initial-boundary value problems, Fokas unified transform method
1 Introduction
The heat equation
$\partial_{t}T-\alpha\Delta T=0,$ (1.1)
where $T$ is temperature, and $\alpha$ is thermal diffusivity, is commonly used for description of heat conduction in homogeneous isotropic media. The heat equation is obtained from the energy balance equation
$C_{\mathrm{v}}\partial_{t}T+\nabla\cdot\boldsymbol{q}=0$ (1.2)
and Fourer’s law (the constitutive equation)
$\boldsymbol{q}=-\kappa\nabla T,$ (1.3)
where $C_{\mathrm{v}}$ is volumetric heat capacity, $\boldsymbol{q}$ is the vector of heat flux, and $\kappa$ is thermal conductivity. Hence the thermal diffusivity is given by $\alpha=\kappa/C_{\mathrm{v}}$.
The energy balance equation (1.2) exactly expresses the first law of thermodynamics as applied to heat transfer: the conservation of thermal energy. Fourier’s law, in turn, is approximate relation between the heat flux and temperature. Fourier’s law can be considered as a phenomenological relation. In classical irreversible (weakly non-equilibrium or linear non-equilibrium) thermodynamics Fourier’s law is derived under the assumption of local thermodynamic equilibrium [1, 2, 3]. However, this assumption is violated in very small dimensions and short timescales, and at low temperatures [4, 5, 6, 7, 8, 9, 10, 11]. Besides, the heat equation violates the causality principle due to infinite speed of heat propagation in this model.
The simplest and apparently most “famous” modification of Fourer’s law is given by the constitutive equation
$\tau\partial_{t}\boldsymbol{q}+\boldsymbol{q}=-\kappa\nabla T,$ (1.4)
where $\tau$ is the relaxation time. This equation is usually attributed to Cattaneo [4] or even to Maxwell, Cattaneo and Vernotte [12, 13]. However, Cattaneo’s priority is questionable [5, 14], see also Ref.[15]. Therefore, we will call this equation modified Fourier’s law. The constitutive equation (1.4) is the first and simplest constitutive equation obtained in the framework of various extended thermodynamics: rational extended thermodynamics [16], extended irreversible thermodynamics [17] and non-equilibrium thermodynamics with internal variables [18], see also Ref. [19].
It follows from the energy balance equation (1.2) and the modified Fourier’s law (1.4) that temperature satisfies the so called hyperbolic heat equation
$\tau\partial_{t}^{2}T+\partial_{t}T-\alpha\Delta T=0.$ (1.5)
Since this equation is hyperbolic, it provides a finite speed of heat propagation and therefore satisfies the causality principle.
Solving initial value problems (in the entire space) for this and other equations obtained in the framework of extended thermodynamics and analysing their solutions is straightforward [20, 21, 22]. However, solving initial-boundary value problems (IBVPs) analytically may encounter difficulties. Below, for simplicity, we will consider equations in one-dimensional space. Also, we will consider problems on a finite interval. Problems on the half line can be solved in the same way.
In one-dimensional space the hyperbolic heat equation (1.5) takes the form
$\tau\partial_{t}^{2}T+\partial_{t}T-\alpha\partial_{x}^{2}T=0.$ (1.6)
If both boundary conditions set the temperature behavior, then the IBVP can easily be solved by the Fourier method or Laplace transform. If both boundary conditions set the heat flux behavior, then the analytical solution for temperature can also be easily found. Indeed, heat flux satisfies the same equation as temperature:
$\tau\partial_{t}^{2}q+\partial_{t}q-\alpha\partial_{x}^{2}q=0.$ (1.7)
First, the IBVP for this equation is solved, and then temperature is determined using the energy balance equation (1.2) (in one dimension).
However, other boundary conditions may be set as well, e. g., one or both boundary conditions can describe Newton’s law (convection boundary conditions), which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case IBVPs cannot be formulated for either equation (1.6) or equation (1.7). In this case IBVPs should be formulated for the system of the energy balance equation (1.2) and the constitutive equation (1.4) (both in one dimension if we stay in one-dimensional space). And in this case, the Fourier method is not applicable. Application of the Laplace transform to these problems is problematic as well. All these problems can be solved by the Fokas unified transform method (UTM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. In this paper we solve these IBVPs by the UTM. The UTM was applied to systems of equations in Refs. [26, 30, 31, 33, 35, 34]. The peculiarity of the problems considered in this paper is that the boundary conditions generally express the relationship of unknown functions (temperature and heat flux) with each other.
In the papers [37, 38, 13, 39, 40] the IBVPs in the framework of the hyperbolic heat equation were solved both analytically and numerically for particular boundary conditions.
2 Statement of the initial-boundary value problem
We consider the IBVP for the system of the energy balance equation (including the heat source term) and modified Fourier’s law on the interval $(0,l)$. For convenience, we consider internal energy $e=C_{\mathrm{v}}T$ instead of temperature $T$. Therefore, we solve the system of the equations
$\partial_{t}e+\partial_{x}q=f(x,t),\quad x\in(0,l),\quad t>0,$ (2.1a)
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0,$ (2.1b)
where $f$ is the heat source term. The system (2.1) is hyperbolic. Initial conditions for the system (2.1) set the initial state of the physical system and are given by
$\left.e\right|_{t=0}=\varphi(x),\quad\left.q\right|_{t=0}=\psi(x).$ (2.2)
The boundary conditions are given by
$\left.(\gamma_{0}e+\delta_{0}q)\right|_{x=0}=g(t),\quad\left.(\gamma_{l}e+\delta_{l}q)\right|_{x=l}=h(t),$ (2.3)
where $|\gamma_{0}|+|\delta_{0}|>0$ and $|\gamma_{l}|+|\delta_{l}|>0$. The general form of the boundary conditions allows us to describe different behavior of the physical system at the boundary: we can set the temperature or heat flux at the boundary; these boundary conditions also include Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. We assume that the initial and boundary conditions are consistent with each other if necessary.
It is convenient to write the system (2.1) in the matrix form
$\partial_{t}\boldsymbol{u}+\varLambda(-\mathrm{i}\partial_{x})\boldsymbol{u}=\boldsymbol{F}(x,t),\quad\boldsymbol{u}=\begin{pmatrix}e\\ q\end{pmatrix},\quad\boldsymbol{F}(x,t)=\begin{pmatrix}f(x,t)\\ 0\end{pmatrix},$ (2.4)
where the matrix $\varLambda$ is given by
$\varLambda(k)=\begin{pmatrix}0&\mathrm{i}k\\[2.0pt] \mathrm{i}k\beta&1/\tau\end{pmatrix},\quad\beta=\frac{\alpha}{\tau}.$
The initial conditions (2.2) take the form
$\left.\boldsymbol{u}\right|_{t=0}=\boldsymbol{u}_{0}(x)\equiv\begin{pmatrix}\varphi(x)\\[2.0pt] \psi(x)\end{pmatrix}.$
3 Solving the problem
3.1 The local relations
First we need to find dispersion relations. It is necessary to note here that the dispersion relations used in the unified transform method differ by a factor of “$\mathrm{i}$” (imaginary unit) from the standard use. We substitute the vector $\boldsymbol{u}=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx-\omega(k)t}\boldsymbol{c}$, where $\boldsymbol{c}$ is an arbitrary constant vector, into the homogeneous system (2.4). Since $\boldsymbol{c}$ is an arbitrary vector, $\omega(k)$ must satisfy the equation
$\det\left[\varLambda(k)-\omega I\right]=0,$
or, equivalently,
$\omega^{2}-\frac{1}{\tau}\omega+\frac{\alpha}{\tau}k^{2}\equiv\omega^{2}-\frac{1}{\tau}\omega+\beta k^{2}=0.$
The roots of the equation, given by
$\omega_{1,2}(k)=\frac{1\pm\sqrt{1-4\alpha\tau k^{2}}}{2\tau},$ (3.1)
determine the dispersion relations.
Next, we need to write the system (2.4) in divergence form, which is called the local relation. We have the chain of equalities
$\partial_{t}\!\left(\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\boldsymbol{u}\right)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\left[\varLambda(k)\boldsymbol{u}+\partial_{t}\boldsymbol{u}\right]\\ =\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\left[\varLambda(k)\boldsymbol{u}-\varLambda(-\mathrm{i}\partial_{x})\boldsymbol{u}+\boldsymbol{F}(x,t)\right]\\ =\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\left[(k+\mathrm{i}\partial_{x})\left.\frac{\varLambda(k)-\varLambda(l)}{k-l}\right|_{l=-\mathrm{i}\partial_{x}}\boldsymbol{u}+\boldsymbol{F}(x,t)\right]\\ =\partial_{x}\!\left(\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}X(k)\boldsymbol{u}\right)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\boldsymbol{F}(x,t),$
where $X$ is a matrix operator, given by
$X(k)=\mathrm{i}\left.\frac{\varLambda(k)-\varLambda(l)}{k-l}\right|_{l=-\mathrm{i}\partial_{x}}.$
So we get the local relation
$\partial_{t}\!\left(\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\boldsymbol{u}\right)-\partial_{x}\!\left(\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}X(k)\boldsymbol{u}\right)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varLambda(k)t}\boldsymbol{F}(x,t).$ (3.2)
In general, the matrix operator $X$ is differential, but this is not the case in this problem. Besides, it does not depend on $k$, and is given by
$X=\begin{pmatrix}0&-1\\ -\beta&0\end{pmatrix}.$
It is convenient to diagonalize the matrix $\varLambda$ (see Refs. [26, 30]):
$\varLambda(k)=S(k)\varOmega(k)S^{-1}(k),$
where $\varOmega$ is a diagonal matrix, given by
$\varOmega(k)=\mathrm{diag}\big{(}\omega_{1}(k),\omega_{2}(k)\big{)},$
so that
$\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}=S(k)\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)t}S^{-1}(k).$
In this case the local relation (3.2) can be written in the form
$\partial_{t}\!\left(\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varOmega(k)t}S^{-1}(k)\boldsymbol{u}\right)-\partial_{x}\!\left(\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varOmega(k)t}S^{-1}(k)X(k)\boldsymbol{u}\right)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{\varOmega(k)t}S^{-1}(k)\boldsymbol{F}(x,t).$ (3.3)
The columns of the matrix $S$ are the eigenvectors of the matrix $\varLambda$, and in this problem
$S(k)=\frac{1}{\omega_{1}-\omega_{2}}\begin{pmatrix}1&1\\ \omega_{1}/\mathrm{i}k&\omega_{2}/\mathrm{i}k\end{pmatrix},\qquad S^{-1}(k)=\begin{pmatrix}-\omega_{2}&\mathrm{i}k\\ \omega_{1}&-\mathrm{i}k\end{pmatrix}.$
3.2 The global relations
Integrating the local relations (3.2) and (3.3) over the domain $\{(x,t)\in(0,l)\times(0,T)\}$, $T>0$, using Green’s theorem, and replacing $T$ with $t$, we obtain the global relations
$\hat{\boldsymbol{u}}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\Lambda(k)t}\hat{\boldsymbol{u}}_{0}(k)+S(k)\boldsymbol{a}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}(k,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\Lambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s,\\ \quad k\in\mathbb{C},$ (3.4)
and
$S^{-1}(k)\hat{\boldsymbol{u}}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)t}S^{-1}(k)\hat{\boldsymbol{u}}_{0}(k)+\boldsymbol{a}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\boldsymbol{b}(k,t)=\tilde{\boldsymbol{F}}(k,t),\quad k\in\mathbb{C},$ (3.5)
respectively, where $\hat{\boldsymbol{u}}$ and $\hat{\boldsymbol{u}}_{0}$ are the Fourier transforms of the solution $\boldsymbol{u}$ and its initial state $\boldsymbol{u}_{0}$, respectively:
$\hat{\boldsymbol{u}}(k,t)=\begin{pmatrix}\hat{e}(k,t)\\ \hat{q}(k,t)\end{pmatrix},\quad\hat{e}(k,t)=\int_{0}^{l}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}e(x,t)\mathop{}\!\mathrm{d}x,\quad\hat{q}(k,t)=\int_{0}^{l}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}q(x,t)\mathop{}\!\mathrm{d}x,$ (3.6)
and
$\hat{\boldsymbol{u}}_{0}(k)=\begin{pmatrix}\hat{\varphi}(k)\\ \hat{\psi}(k)\end{pmatrix},\quad\hat{\varphi}(k)=\int_{0}^{l}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\varphi(x)\mathop{}\!\mathrm{d}x,\quad\hat{\psi}(k)=\int_{0}^{l}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}\psi(x)\mathop{}\!\mathrm{d}x,$ (3.7)
(we assume that all the functions outside the interval $[0,l]$ are zero, therefore, $\hat{}$ means the Fourier transform), the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ are expressed through the boundary values of the solution $\boldsymbol{u}$ by
$\boldsymbol{a}(k,t)$ $=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)X\boldsymbol{u}(0,s)\mathop{}\!\mathrm{d}s=\begin{pmatrix}-\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{q}^{(1)}(0,t)\\[3.0pt] \mathrm{i}k\beta\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{q}^{(2)}(0,t)\end{pmatrix},$ (3.8a)
$\boldsymbol{b}(k,t)$ $=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)X\boldsymbol{u}(l,s)\mathop{}\!\mathrm{d}s=\begin{pmatrix}-\mathrm{i}k\beta\tilde{e}^{(1)}(l,t)+\omega_{2}\tilde{q}^{(1)}(l,t)\\[3.0pt] \mathrm{i}k\beta\tilde{e}^{(2)}(l,t)-\omega_{1}\tilde{q}^{(2)}(l,t)\end{pmatrix},$ (3.8b)
the functions $\tilde{e}^{(m)}$ and $\tilde{q}^{(m)}$ are time transforms of the boundary values of $e$ and $q$, respectively, defined by [^1] This time transform differs from the time transform, used in the UTM literature. However, this form of the transform seems to be more convenient.
$\tilde{e}^{(m)}(x,t)\equiv\tilde{e}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}e(x,s)\mathop{}\!\mathrm{d}s,$ (3.9a)
$\tilde{q}^{(m)}(x,t)\equiv\tilde{q}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}q(x,s)\mathop{}\!\mathrm{d}s,$ (3.9b)
(note that $\tilde{e}^{(m)}$ and $\tilde{q}^{(m)}$ do not change under the transformation $k\to-k$), $\tilde{\boldsymbol{F}}$ is expressed through the Fourier transform of $\boldsymbol{F}$ by
$\tilde{\boldsymbol{F}}(k,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s,$
$\hat{\boldsymbol{F}}(k,t)=\begin{pmatrix}\hat{f}(k,t)\\ 0\end{pmatrix},\quad\hat{f}(k,t)=\int_{0}^{l}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kx}f(x,t)\mathop{}\!\mathrm{d}x,$ (3.10)
$\hat{f}$ is the Fourier transform of $f$. $\tilde{\boldsymbol{F}}$ can be easily expressed through the Fourier transform of $f$:
$\tilde{\boldsymbol{F}}(k,t)\equiv\begin{pmatrix}\tilde{F}_{1}(k,t)\\[2.0pt] \tilde{F}_{2}(k,t)\end{pmatrix}=\int_{0}^{t}\begin{pmatrix}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}(t-s)}\\[2.0pt] \omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}(t-s)}\end{pmatrix}\hat{f}(k,s)\mathop{}\!\mathrm{d}s.$ (3.11)
It is important to note that the global relations (3.4) and (3.5) are valid for $k\in\mathbb{C}$. Indeed, the spatial variable $x$ belongs to the finite interval $[0,l]$, therefore, if we assume sufficient decay of the functions $e$, $q$, $\varphi$, $\psi$ and $f$ as $x\to\infty$ for all $t$, then the Fourier transforms can be analytically continued from the real axis of the spectral parameter $k$ onto the entire complex $k$-plane due to the exponential decay there.
3.3 A preliminary expression for the solution
We assume that $e$, $q$, $\varphi$, $\psi$ and $f$ are zero, when $x<0$ or $x>l$. Therefore, the transform $\hat{}$ in Eqs. (3.6), (3.7) and (3.10) is the Fourier transform over the entire real axis. Applying the inverse Fourier transform to the global relation (3.4), we obtain the preliminary integral expression for the solution
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}(k,t)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}k(l-x)}S(k)\boldsymbol{b}(k,t)\mathop{}\!\mathrm{d}k.$ (3.12)
This formula is not yet the final representation of the solution to the problem (2.1)–(2.3). Indeed, the formula contains the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The vectors depend on the functions $\tilde{e}^{(1,2)}$ and $\tilde{q}^{(1,2)}$ (see Eq. (3.8)) that are time transforms of the boundary values of the unknown functions $e$ and $q$ (see Eq. (3.9)). Depending on the coefficients in the boundary conditions (2.3), some of these boundary functions or all of them are unknown and need to be found.
3.4 Deformation of the integration path
In order to obtain a genuine representation of the solution, we first need to deform the integration path in the integrals involving the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ in Eq. (3.12) from the real axis to appropriate paths in the complex $k$-plane. We define the domains $D^{\pm}=D\cap\mathbb{C}^{\pm}$, where ±=$\mathbb{C}^{\pm}=\{k\in\mathbb{C}:\operatorname{Im}k\gtrless 0\}$, and
$D=\{k\in\mathbb{C}:\operatorname{Re}\omega_{1}(k)<0\text{ or }\operatorname{Re}\omega_{2}(k)<0\}\equiv\left\{k=k_{\mathrm{R}}+\mathrm{i}k_{\mathrm{I}}:k_{\mathrm{I}}^{2}\geq\frac{k_{\mathrm{R}}^{2}}{4\alpha\tau k_{\mathrm{R}}^{2}+1}\right\}.$
Using the same reasoning as in Refs. [24, 25, 27, 29, 30], we transform the formula (3.12) into the formula
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}k(l-x)}S(k)\boldsymbol{b}(k,t)\mathop{}\!\mathrm{d}k,$ (3.13)
where
$\partial D^{\pm}=\left\{k=k_{\mathrm{R}}+\mathrm{i}k_{\mathrm{I}}:k_{\mathrm{I}}=\pm\sqrt{\frac{k_{\mathrm{R}}^{2}}{4\alpha\tau k_{\mathrm{R}}^{2}+1}},\ k_{\mathrm{R}}\in(-\infty,\infty)\right\},$
and we assume that the path of integration along the boundary of a domain is directed in such a way that the domain is to the left of the boundary when it is traversed. Besides, we deform the paths of integration from $\partial D^{\pm}$ to $\partial D_{0}^{\pm}$ in the vicinity of the origin. Note that the imaginary part of the spectral parameter $k$ on the integration paths is bounded.
3.5 Using the symmetry of the dispersion relations
Next, we need to use the symmetry of the dispersion relations. They are invariant under the transformation $k\to-k$: $\omega_{1,2}(-k)=\omega_{1,2}(k)$. Using this transformation in the global relation (3.5) we obtain another global relation (note that $\varOmega(-k)=\varOmega(k)$)
$S^{-1}(-k)\hat{\boldsymbol{u}}(-k,t)-\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)t}S^{-1}(-k)\hat{\boldsymbol{u}}_{0}(-k)+\boldsymbol{a}(-k,t)-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\boldsymbol{b}(-k,t)\\ =\tilde{\boldsymbol{F}}(-k,t),\quad k\in\mathbb{C},$ (3.14)
To make formulae more compact, we define the vectors
$\boldsymbol{U}(k,t)\equiv\begin{pmatrix}U_{1}(k,t)\\[2.0pt] U_{2}(k,t)\end{pmatrix}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}S^{-1}(k)\hat{\boldsymbol{u}}(k,t)=\begin{pmatrix}-\omega_{2}\\[2.0pt] \omega_{1}\end{pmatrix}\hat{e}(k,t)+\begin{pmatrix}\mathrm{i}k\\[2.0pt] -\mathrm{i}k\end{pmatrix}\hat{q}(k,t),$ (3.15)
$\boldsymbol{\varPhi}(k)\equiv\begin{pmatrix}\varPhi_{1}(k)\\[2.0pt] \varPhi_{2}(k)\end{pmatrix}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)t}S^{-1}(k)\hat{\boldsymbol{u}}_{0}(k)=\begin{pmatrix}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}\\[2.0pt] \omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}\end{pmatrix}\hat{\varphi}(k)+\begin{pmatrix}\mathrm{i}k\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}\\[2.0pt] -\mathrm{i}k\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}\end{pmatrix}\hat{\psi}(k).$ (3.16)
Then the global relations (3.5), (3.14) can be written in the form
$\boldsymbol{a}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\boldsymbol{b}(k,t)$ $=-\boldsymbol{U}(k,t)+\boldsymbol{\varPhi}(k,t)+\tilde{\boldsymbol{F}}(k,t),\quad k\in\mathbb{C},$ (3.17a)
$\boldsymbol{a}(-k,t)-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\boldsymbol{b}(-k,t)$ $=-\boldsymbol{U}(-k,t)+\boldsymbol{\varPhi}(-k,t)+\tilde{\boldsymbol{F}}(-k,t),\quad k\in\mathbb{C},$ (3.17b)
respectively, where $\boldsymbol{a}$ and $\boldsymbol{b}$ are given by Eq. (3.8), and $\tilde{\boldsymbol{F}}$ is given by Eq. (3.11).
Now we apply the time transform to the boundary conditions (2.3) and obtain the relations
$\gamma_{0}\tilde{e}^{(m)}(0,t)+\delta_{0}\tilde{q}^{(m)}(0,t)=\tilde{g}^{(m)}(t),\quad\gamma_{l}\tilde{e}^{(m)}(l,t)+\delta_{l}\tilde{q}^{(m)}(l,t)=\tilde{h}^{(m)}(t),$ (3.18)
where
$\tilde{g}^{(m)}(t)\equiv\tilde{g}(\omega_{m}(k),t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}g(s)\mathop{}\!\mathrm{d}s,$ (3.19a)
$\tilde{h}^{(m)}(t)\equiv\tilde{h}(\omega_{m}(k),t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}h(s)\mathop{}\!\mathrm{d}s.$ (3.19b)
Note that $\tilde{g}^{(m)}$ and $\tilde{h}^{(m)}$ do not change under the transformation $k\to-k$.
To avoid too cumbersome general calculations, below we will limit ourselves to the most interesting boundary conditions from a physical point of view. We will limit ourselves to boundary conditions that determine the heat flux or express Newton’s law. We will not consider the conditions determining the boundary temperature. Therefore, we assume that $\delta_{0}=1$, $\delta_{l}=-1$ and $\gamma_{0,l}\geq 0$. In this case the boundary conditions (2.3) take the form
$\left.(\gamma_{0}e+q)\right|_{x=0}=g(t),\quad\left.(\gamma_{l}e-q)\right|_{x=l}=h(t),$ (3.20)
and the relations (3.18) take the form
$\tilde{q}^{(m)}(0,t)=\tilde{g}^{(m)}(t)-\gamma_{0}\tilde{e}^{(m)}(0,t),\quad\tilde{q}^{(m)}(l,t)=\gamma_{l}\tilde{e}^{(m)}(l,t)-\tilde{h}^{(m)}(t).$ (3.21)
Using the relations (3.21) we can express the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ (Eq. (3.8)) via $\tilde{e}^{(m)}$, $\tilde{g}^{(m)}$ and $\tilde{h}^{(m)}$:
$\boldsymbol{a}(k,t)=\begin{pmatrix}-\sigma_{0}^{(2)}(k)\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)\\[4.0pt] \sigma_{0}^{(1)}(k)\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{g}^{(2)}(t)\end{pmatrix},$ (3.22a)
$\boldsymbol{b}(k,t)=\begin{pmatrix}\sigma_{l}^{(2)}(-k)\tilde{e}^{(1)}(l,t)-\omega_{2}\tilde{h}^{(1)}(t)\\[4.0pt] -\sigma_{l}^{(1)}(-k)\tilde{e}^{(2)}(l,t)+\omega_{1}\tilde{h}^{(2)}(t)\end{pmatrix},$ (3.22b)
where
$\sigma_{0,l}^{(m)}(k)=\mathrm{i}k\beta+\gamma_{0,l}\omega_{m}(k).$ (3.23)
The vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ depend on the functions $\tilde{e}^{(1,2)}$, which are still unknown. To find them, we use the system (3.17) of the global relations and the representations (3.22). The system (3.17), consisting of four equations, can be rewritten as the system of two algebraic systems
$\begin{pmatrix}\sigma_{0}^{(2)}(k)&\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[4.0pt] \sigma_{0}^{(2)}(-k)&\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\end{pmatrix}\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[4.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\begin{pmatrix}U_{1}(k,t)-V_{1}(k,t)\\[4.0pt] U_{1}(-k,t)-V_{1}(-k,t)\end{pmatrix},$ (3.24a)
$\begin{pmatrix}\sigma_{0}^{(1)}(k)&\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[2.0pt] \sigma_{0}^{(1)}(-k)&\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\end{pmatrix}\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\[2.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=-\begin{pmatrix}U_{2}(k,t)-V_{2}(k,t)\\[2.0pt] U_{2}(-k,t)-V_{2}(-k,t)\end{pmatrix}$ (3.24b)
(the equations of the system (3.24a) are the first equations (raws) of the systems (3.17a) and (3.17b), and the equations of the system (3.24b) are the second equations (raws) of these systems), where
$\begin{split}V_{1}(k,t)=\varPhi_{1}(k)+\tilde{F}_{1}(k,t)-\omega_{2}\left[\tilde{g}^{(1)}(t)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\tilde{h}^{(1)}(t)\right],\\ V_{2}(k,t)=\varPhi_{2}(k)+\tilde{F}_{2}(k,t)+\omega_{1}\left[\tilde{g}^{(2)}(t)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\tilde{h}^{(2)}(t)\right],\end{split}$ (3.25)
$\varPhi_{1,2}$ are given by Eq. (3.16), $\tilde{F}_{1,2}$ are given by Eq. (3.11), $\tilde{g}^{(1,2)}$ and $\tilde{h}^{(1,2)}$ are given by Eq. (3.19). The inverse of the matrices in the systems (3.24a) and (3.24b) are given by
$\frac{1}{\varDelta_{1}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)&-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[4.0pt] -\sigma_{0}^{(2)}(-k)&\sigma_{0}^{(2)}(k)\end{pmatrix}\quad\text{and}\quad\frac{1}{\varDelta_{2}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)&-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[4.0pt] -\sigma_{0}^{(1)}(-k)&\sigma_{0}^{(1)}(k)\end{pmatrix},$
respectively, where
$\begin{split}\varDelta_{1}(k)&=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{0}^{(2)}(k)\sigma_{l}^{(2)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{0}^{(2)}(-k)\sigma_{l}^{(2)}(-k),\\[2.0pt] \varDelta_{2}(k)&=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{0}^{(1)}(k)\sigma_{l}^{(1)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{0}^{(1)}(-k)\sigma_{l}^{(1)}(-k).\end{split}$ (3.26)
Solving the systems (3.24), we find the functions $\tilde{e}^{(1,2)}$:
$\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[2.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}=\frac{1}{\varDelta_{1}(k)}\Bigg{[}$ $\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\\[2.0pt] -\sigma_{0}^{(2)}(-k)\end{pmatrix}\left[U_{1}(k,t)-V_{1}(k,t)\right]$ (3.27a)
$+\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[2.0pt] \sigma_{0}^{(2)}(k)\end{pmatrix}\left[U_{1}(-k,t)-V_{1}(-k,t)\right]\Bigg{]},$
$\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\[2.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}=\frac{1}{\varDelta_{2}(k)}\Bigg{[}$ $\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\\[2.0pt] \sigma_{0}^{(1)}(-k)\end{pmatrix}\left[U_{2}(k,t)-V_{2}(k,t)\right]$ (3.27b)
$+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[2.0pt] -\sigma_{0}^{(1)}(k)\end{pmatrix}\left[U_{2}(-k,t)-V_{2}(-k,t)\right]\Bigg{]},$
where $U_{1,2}$ are given by Eq. (3.15) and $V_{1,2}$ are given by Eq. (3.25). Note that the functions $U_{1,2}$ are still unknown, while the functions $V_{1,2}$ are known, they contain only known functions.
3.6 The solution
The formula (3.13) is also not the final representation of the solution, because the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ depend on the unknown functions $e$ and $q$ through the functions $U_{1,2}$ (see Eqs. (3.22), (3.27) and (3.15)). However, it can be shown that integrals containing the functions $U_{1,2}$ are equal to zero, see, e. g., Refs. [25, 27, 29, 30], and, therefore, the final formula does not contain these integrals.
As a result, the solution is given by
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}k(l-x)}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (3.28)
where $\hat{\boldsymbol{u}}_{0}(k)$ is given by Eq. (3.7), $\hat{\boldsymbol{F}}(k,t)$ is given by Eq. (3.10),
$\boldsymbol{a}^{\prime}(k,t)=\begin{pmatrix}-\sigma_{0}^{(2)}(k)\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)\\[4.0pt] \sigma_{0}^{(1)}(k)\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{g}^{(2)}(t)\end{pmatrix},$ (3.29a)
$\boldsymbol{b}^{\prime}(k,t)=\begin{pmatrix}\sigma_{l}^{(2)}(-k)\tilde{e}^{(1)}(l,t)-\omega_{2}\tilde{h}^{(1)}(t)\\[4.0pt] -\sigma_{l}^{(1)}(-k)\tilde{e}^{(2)}(l,t)+\omega_{1}\tilde{h}^{(2)}(t)\end{pmatrix},$ (3.29b)
(the expressions for $\boldsymbol{a}^{\prime}$ and $\boldsymbol{b}^{\prime}$ formally coincide with the expressions for $\boldsymbol{a}$ and $\boldsymbol{b}$, see Eq. (3.22), however, the functions $\tilde{e}^{(1,2)}$ are different, see Eq. (3.30) below), the coefficients $\sigma_{0,l}^{(1,2)}$ are given by Eq. (3.23), the functions $\tilde{g}^{(1,2)}$ and $\tilde{h}^{(1,2)}$ are given by Eq. (3.19), the functions $\tilde{e}^{(1,2)}$ are given by
$\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[3.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{1}(k)}\Bigg{[}\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)\\[3.0pt] \sigma_{0}^{(2)}(-k)\end{pmatrix}V_{1}(k,t)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[3.0pt] -\sigma_{0}^{(2)}(k)\end{pmatrix}V_{1}(-k,t)\Bigg{]},$ (3.30a)
$\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\[3.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{2}(k)}\Bigg{[}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)\\[3.0pt] -\sigma_{0}^{(1)}(-k)\end{pmatrix}V_{2}(k,t)+\begin{pmatrix}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[3.0pt] \sigma_{0}^{(1)}(k)\end{pmatrix}V_{2}(-k,t)\Bigg{]}$ (3.30b)
(these are Eqs. (3.27) with $U_{1,2}=0$), the functions $V_{1,2}$ are given by Eq. (3.25), the coefficients $\varDelta_{1,2}(k)$ are given by Eq. (3.26).
4 A numerical example
To illustrate the representation (3.28) of the solution, we consider here a particular case, in which heat sources are absent ($f=0$), the initial state is zero ($\varphi=0$, $\psi=0$), the heat flux is prescribed at the left boundary ($\gamma_{0}=0$), and heat exchange at the right boundary obeys Newton’s law, while the ambient temperature is zero ($h=0$), In this case both equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (4.1a)
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (4.1b)
the initial conditions are also homogeneous:
$\left.e\right|_{t=0}=0,\quad\left.q\right|_{t=0}=0,$ (4.2)
and the boundary conditions take the form
$\left.q\right|_{x=0}=g(t),\quad\left.q\right|_{x=l}=\left.\gamma_{l}e\right|_{x=l}.$ (4.3)
The solution to the problem (4.1)–(4.3) is given by
$\boldsymbol{u}(x,t)=-\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (4.4)
where the functions $\boldsymbol{a}^{\prime}$ and $\boldsymbol{b}^{\prime}$ are given by
$\boldsymbol{a}^{\prime}(k,t)$ $=\begin{pmatrix}-\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)\\[4.0pt] \mathrm{i}k\beta\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{g}^{(2)}(t)\end{pmatrix},$
$\boldsymbol{b}^{\prime}(k,t)$ $=\begin{pmatrix}\sigma_{l}^{(2)}(-k)\tilde{e}^{(1)}(l,t)\\[4.0pt] -\sigma_{l}^{(1)}(-k)\tilde{e}^{(2)}(l,t)\end{pmatrix},$
the functions $\tilde{e}^{(1,2)}$ are given by
$\begin{pmatrix}\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)\\[3.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{1}^{\prime}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[3.0pt] 2\end{pmatrix}\omega_{2}\tilde{g}^{(1)}(t),$ (4.5a)
$\begin{pmatrix}\mathrm{i}k\beta\tilde{e}^{(2)}(0,t)\\[3.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{2}^{\prime}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[3.0pt] 2\end{pmatrix}\omega_{1}\tilde{g}^{(2)}(t),$ (4.5b)
the coefficients $\varDelta_{1,2}^{\prime}(k)$ are given by
$\varDelta_{1}^{\prime}(k)=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k),$
$\varDelta_{2}^{\prime}(k)=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k).$
Using the relations (4.5), we see that
$-\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)=2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\frac{\sigma_{l}^{(2)}(-k)}{\varDelta_{1}^{\prime}(k)}\omega_{2}\tilde{g}^{(1)}(t),\quad\tilde{e}^{(1)}(l,t)=\frac{2}{\varDelta_{1}^{\prime}(k)}\omega_{2}\tilde{g}^{(1)}(t),$
$-\mathrm{i}k\beta\tilde{e}^{(2)}(0,t)+\omega_{1}\tilde{g}^{(2)}(t)=2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\frac{\sigma_{l}^{(1)}(-k)}{\varDelta_{2}^{\prime}(k)}\omega_{1}\tilde{g}^{(2)}(t),\quad\tilde{e}^{(2)}(l,t)=\frac{2}{\varDelta_{2}^{\prime}(k)}\omega_{1}\tilde{g}^{(2)}(t).$
Hence
$S(k)\boldsymbol{a}^{\prime}(k,t)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\\[2.0pt] =\frac{2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}{\omega_{1}-\omega_{2}}\bigg{[}\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\frac{\sigma_{l}^{(2)}(-k)}{\varDelta_{1}^{\prime}(k)}\omega_{2}\tilde{g}^{(1)}(t)-\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\frac{\sigma_{l}^{(1)}(-k)}{\varDelta_{2}^{\prime}(k)}\omega_{1}\tilde{g}^{(2)}(t)\bigg{]}.$ (4.6)
As a result, we conclude that the solution to the problem (4.1)–(4.3) is given by the formula (4.4), where $S\boldsymbol{a}^{\prime}$ and $S\boldsymbol{b}^{\prime}$ are given by Eq. (4.6).
As a function, determining the heat flux on the left boundary, we choose the function
$g(t)=\frac{2}{\tau_{\Delta}}\sin^{2}\!\left(\frac{\pi t}{\tau_{\Delta}}\right)\mathbbm{1}_{(0,\tau_{\Delta})}(t),$ (4.7)
where
$\mathbbm{1}_{A}(t)=\begin{cases}1,&t\in A,\\ 0,&t\notin A,\end{cases}$
is the indicator function of a set $A$. Note that $\int_{0}^{\infty}g(t)\mathop{}\!\mathrm{d}t\equiv\int_{0}^{\tau_{\Delta}}g(t)\mathop{}\!\mathrm{d}t=1$. The problem (4.1)–(4.3) with the heat flux of the form (4.7) simulates laser flash experiments [37].
Fig. 1 shows solutions to the problem (4.1)–(4.3) obtained by using the formula (4.4). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$. Here we use dimensionless values. The solution is calculated on the right boundary, i. e., for $x=1$. The solution to the problem with $\gamma_{l}=0$ (the thermally insulating boundary) and the general heat flux $g(t)$ can easily be obtained by the Fourier method with preliminary elimination of inhomogeneities in boundary conditions. This solution is given by Eq. (A.5) (we have checked that in this case the formulae (4.4) and (A.5) give the same numerical results). In Ref. [37] the series solution to the problem (4.1)–(4.3) with $\gamma_{l}=0$ was obtained for the heat flux of the form (4.7).
Figure 1: The solutions to the problem (4.1)–(4.3) on the right boundary, i. e., for $x=1$ with the heat flux $g(t)$ given by Eq. (4.7). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$.
5 Conclusions
In this paper we have considered the IBVPs on the finite interval for the system of the energy balance equation and modified Fourier’s law (constitutive equation). This system is hyperbolic. The boundary conditions include various models of behavior of a physical system at the boundaries. The most interesting model is when the behavior on at least one of the boundaries is described by Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case the boundary conditions express the relationship of unknown functions (temperature or internal energy and heat flux) with each other.
To solve the problems, we have applied the unified transform method. We did not solve problems in which one of the boundary conditions determines the temperature. These problems can be solved in the same way as in this paper. To illustrate the final formulae, we have considered a numerical example for a special case in which heat exchange at one of the boundaries obeys Newton’s law.
Also of considerable interest is the integral representation of solutions to IBVPs in the framework of the Guyer–Krumhansl model of heat conduction [41, 4, 5, 17, 42]. In the papers [37, 38, 43, 44, 45, 40] the IBVPs in the framework of the Guyer–Krumhansl model were solved both analytically and numerically for particular boundary conditions. The integral representation of the solutions to the initial value problems in the framework of the Guyer–Krumhansl model will be considered elsewhere.
Appendix A Particular cases that allow series representation of the solutions
In this appendix, for simplicity, we assume that heat sources are absent, i. e., $f=0$.
A.1 The initial state is zero, the heat flux is prescribed on the left boundary, and the right boundary is thermally insulating
Here we assume that the initial state is zero ($\varphi=0$, $\psi=0$), the heat flux is prescribed on the left boundary ($\gamma_{0}=0$), and the right boundary is thermally insulating ($\gamma_{l}=0$ and $h=0$). In this case both equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (A.1a)
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (A.1b)
the initial conditions are also homogeneous:
$\left.e\right|_{t=0}=0,\quad\left.q\right|_{t=0}=0,$ (A.2)
and the boundary conditions take the form
$\left.q\right|_{x=0}=g(t),\quad\left.q\right|_{x=l}=0.$ (A.3)
This is a particular case of the problem (4.1)–(4.3), where $\gamma_{l}=0$.
In this case the solution to the problem (A.1)–(A.3) is given by
$\boldsymbol{u}(x,t)=-\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (A.4)
where
$S(k)\boldsymbol{a}^{\prime}(k,t)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\\[2.0pt] =\frac{2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}{(\omega_{1}-\omega_{2})(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})}\bigg{[}\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\omega_{1}\tilde{g}^{(2)}(t)-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\omega_{2}\tilde{g}^{(1)}(t)\bigg{]}.$
In order to obtain a series representation, the paths of integration $\partial D_{0}^{\pm}$ must be deformed back to the real axis. Note that the branch points $k=\pm(2\sqrt{\alpha\tau})^{-1}$ are removable. By deforming the paths, avoiding poles located on the real axis, and taking into account opposite directions of integration, we obtain contours around the poles. As a result of integration along the contours, the representation (A.4) is transformed into the series
$\boldsymbol{u}(x,t)=\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right],$
where $k_{n}=\pi n/l$ are the poles.
Using the equalities $\mathop{\mathrm{e}}\nolimits^{\pm\mathrm{i}k_{n}l}=(-1)^{n}$, $[(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})_{k}^{\prime}]|_{k=k_{n}}=2\mathrm{i}l(-1)^{n}$ and $\omega_{1}(k_{n})\omega_{2}(k_{n})=\beta k_{n}^{2}$, we see that the residues are equal to
$\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{\mathrm{i}l(\omega_{1}^{(n)}-\omega_{2}^{(n)})}\begin{pmatrix}\omega_{1}^{(n)}\tilde{g}^{(2)}(t)-\omega_{2}^{(n)}\tilde{g}^{(1)}(t)\\[5.0pt] -\mathrm{i}k_{n}\beta\left[\tilde{g}^{(2)}(t)-\tilde{g}^{(1)}(t)\right]\end{pmatrix},\quad n\neq 0,$
and
$\operatorname*{Res}_{k=0}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=0}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]=\frac{1}{\mathrm{i}l}\begin{pmatrix}\tilde{g}\\[5.0pt] 0\end{pmatrix},$
where $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$, and
$\tilde{g}(t)=\int_{0}^{t}g(s)\mathop{}\!\mathrm{d}s.$
As a result, we conclude that the solution is given by
$\boldsymbol{u}(x,t)\equiv\begin{pmatrix}e(x,t)\\[2.0pt] q(x,t)\end{pmatrix}\\[2.0pt] =\frac{1}{l}\begin{pmatrix}\tilde{g}(t)\\[3.0pt] 0\end{pmatrix}+\frac{2}{l}\sum_{n=1}^{\infty}\frac{1}{\omega_{1}^{(n)}-\omega_{2}^{(n)}}\begin{pmatrix}\left[\omega_{1}^{(n)}\tilde{g}^{(2)}(t)-\omega_{2}^{(n)}\tilde{g}^{(1)}(t)\right]\cos(k_{n}x)\\[3.0pt] k_{n}\beta\left[\tilde{g}^{(2)}(t)-\tilde{g}^{(1)}(t)\right]\sin(k_{n}x)\end{pmatrix},$ (A.5)
where $k_{n}=\pi n/l$, $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$ (see Eq. (3.1)), $\tilde{g}^{(1,2)}(t)$ are given by Eq. (3.19a). This solution can also be easily obtained by the Fourier method with preliminary elimination of inhomogeneities in boundary conditions. In Ref. [37] the solution was obtained for a particular heat flux $g(t)$ of the form (4.7).
A.2 The initial state is non-zero and the boundaries are thermally insulating
Here we assume that the initial state is non-zero and the boundaries are thermally insulating ($\gamma_{0}=\gamma_{l}=0$ and $g=0$, $h=0$). In this case both equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (A.6a)
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (A.6b)
the initial conditions remain the same:
$\left.e\right|_{t=0}=\varphi(x),\quad\left.q\right|_{t=0}=\psi(x),$ (A.7)
and the boundary conditions take the form
$\left.q\right|_{x=0}=0,\quad\left.q\right|_{x=l}=0.$ (A.8)
The solution to the problem (A.6)–(A.8) is given by
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (A.9)
where $\hat{\boldsymbol{u}}_{0}(k)$ is given by Eq. (3.7),
$S(k)\boldsymbol{a}^{\prime}(k,t)$ $=\frac{\mathrm{i}k\beta}{\omega_{1}-\omega_{2}}\bigg{[}-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(1)}(0,t)+\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(2)}(0,t)\bigg{]},$ (A.10a)
$S(k)\boldsymbol{b}^{\prime}(k,t)$ $=\frac{\mathrm{i}k\beta}{\omega_{1}-\omega_{2}}\bigg{[}-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(1)}(l,t)+\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(2)}(l,t)\bigg{]},$ (A.10b)
(see Eq. (3.29), where $\tilde{g}^{(1,2)}=0$, $\tilde{h}^{(1,2)}=0$ and $\sigma_{0,l}^{(1,2)}(k)=\mathrm{i}k\beta$), the functions $\tilde{e}^{(1,2)}$ are given by
$-\mathrm{i}k\beta\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[2.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}\bigg{[}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{1}(k)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{1}(-k)\bigg{]},$ (A.11a)
$\mathrm{i}k\beta\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\ \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}\bigg{[}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{2}(k)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{2}(-k)\bigg{]},$ (A.11b)
(see Eq. (3.30), where $V_{1,2}(k,t)=\varPhi_{1,2}(k)$ and $\varDelta_{1,2}(k)=-(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})k^{2}\beta^{2}$).
Deforming the paths $\partial D_{0}^{\pm}$ back to the real axis, we see that the representation (A.9) is transformed into the representation
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]\\ +\frac{1}{2\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right],$ (A.12)
where $k_{n}=\pi n/l$ are the poles. It follows from Eqs. (A.10) and (A.11) the relation
$S(k)\boldsymbol{a}^{\prime}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)=\frac{1}{\omega_{1}-\omega_{2}}\begin{pmatrix}\varPhi_{1}(k)+\varPhi_{2}(k)\\[2.0pt] \left[\omega_{1}\varPhi_{1}(k)+\omega_{2}\varPhi_{2}(k)\right]/\mathrm{i}k\end{pmatrix}\\[2.0pt] =\frac{1}{\omega_{1}-\omega_{2}}\left[\begin{pmatrix}(\omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t})\\[2.0pt] \mathrm{i}k\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t})\end{pmatrix}\hat{\varphi}(k)+\begin{pmatrix}\mathrm{i}k(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t})\\[2.0pt] (\omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t})\end{pmatrix}\hat{\psi}(k)\right]\\[2.0pt] =\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)$
Hence the difference of the two principal-value integrals in Eq. (A.12) is equal to the first integral, and, therefore, the net contribution of the three integrals to the solution is equal to zero. Thus the representation (A.12) takes the form
$\boldsymbol{u}(x,t)=\frac{\mathrm{i}}{2}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]+\frac{\mathrm{i}}{2}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right].$
Using the equalities $\mathop{\mathrm{e}}\nolimits^{\pm\mathrm{i}k_{n}l}=(-1)^{n}$ and $[(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})_{k}^{\prime}]|_{k=k_{n}}=2\mathrm{i}l(-1)^{n}$ we see that the residues are equal to
$\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{2\mathrm{i}l(\omega_{1}^{(n)}-\omega_{2}^{(n)})}\begin{pmatrix}\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})+\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\\[2.0pt] \left\{\omega_{1}^{(n)}\left[\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})\right]+\omega_{2}^{(n)}\left[\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\right]\right\}/\mathrm{i}k_{n}\end{pmatrix},$
where $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$. The functions $\varPhi_{1,2}$ are given by Eq. (3.16), hence
$\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})+\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\\[2.0pt] =(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t})\left[\hat{\varphi}(k_{n})+\hat{\varphi}(-k_{n})\right]\\[2.0pt] +\mathrm{i}k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\left[\hat{\psi}(k_{n})-\hat{\psi}(-k_{n})\right]\\[2.0pt] =2\left[(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t})\varphi_{n}+k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\right],\quad n\in\mathbb{Z}\backslash\{0\},$
and
$\frac{\omega_{1}^{(n)}\left[\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})\right]+\omega_{2}^{(n)}\left[\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\right]}{\mathrm{i}k_{n}}\\[2.0pt] =\mathrm{i}k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\left[\hat{\varphi}(k_{n})+\hat{\varphi}(-k_{n})\right]\\[2.0pt] +(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\left[\hat{\psi}(k_{n})-\hat{\psi}(-k_{n})\right]\\[2.0pt] =2\mathrm{i}\left[k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\varphi_{n}-(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\right],\quad n\in\mathbb{Z}\backslash\{0\},$
where
$\hat{\varphi}(k_{n})+\hat{\varphi}(-k_{n})=l\varphi_{n},\quad\varphi_{n}=\frac{2}{l}\int_{0}^{l}\varphi(x)\cos(k_{n}x)\mathop{}\!\mathrm{d}x,\quad\varphi_{0}=\frac{1}{l}\int_{0}^{l}\varphi(x)\mathop{}\!\mathrm{d}x,$
$\hat{\psi}(k_{n})-\hat{\psi}(-k_{n})=-\mathrm{i}l\psi_{n},\quad\psi_{n}=\frac{2}{l}\int_{0}^{l}\psi(x)\sin(k_{n}x)\mathop{}\!\mathrm{d}x.$
As a result, we conclude that the solution is given by
$\boldsymbol{u}(x,t)\equiv\begin{pmatrix}e(x,t)\\[2.0pt] q(x,t)\end{pmatrix}=\begin{pmatrix}\varphi_{0}\\[3.0pt] 0\end{pmatrix}+\sum_{n=1}^{\infty}\frac{1}{\omega_{1}^{(n)}-\omega_{2}^{(n)}}\times\\[4.0pt] \times\begin{pmatrix}\big{[}(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t})\varphi_{n}+k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\big{]}\cos(k_{n}x)\\[4.0pt] \big{[}-k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\varphi_{n}+(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\big{]}\sin(k_{n}x)\end{pmatrix},$
where $k_{n}=\pi n/l$, $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$ (see Eq. (3.1)). This solution can also be easily obtained by the Fourier method (see the solution in Ref. [13]).